# The theory of implicit operations

### v1

@@ 10. Primitive positive expansions
Let $f$ be the unary implicit operation of the quasivariety $\mathsf{RCRing}$ of reduced commutative rings given by Theorem 8.18 and recall from the same theorem that $f\in\mathsf{ext}_{pp}(\mathsf{RCRing})$. Moreover, let $(\,)^{*}$ be a unary function symbol. Then the language ${L}_{f}={L}_{\mathsf{RCRing}}\cup\{(\,)^{*}\}$ is an $f$-expansion of ${L}_{\mathsf{RCRing}}$, in which the role of $g_{f}$ is played by $(\,)^{*}$.
@@
where $\mathsf{Meadow}$ is the variety of meadows. Lastly, recall that $f\in\mathsf{ext}_{pp}(\mathsf{RCRing})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{Field}$. Together with the left hand side of the above display, this allows us to apply 10.5, obtaining that $\mathbb{Q}(\mathsf{Field}[{L}_{f}])$ is a pp expansion of $\mathsf{RCRing}$. By applying in succession the equality $\mathsf{Field}[{L}_{f}]=\mathsf{Field}^{*}$ and the right hand side of the above display, we obtain
@@
Now, let $f\in\mathsf{ext}_{pp}(\mathsf{Hilbert})$ be the binary implicit operation given by Theorem 8.22. Moreover, let $\wedge$ be a binary function symbol. Then the language ${L}_{f}={L}_{\mathsf{Hilbert}}\cup\{\wedge\}$ is an $f$-expansion of ${L}_{\mathsf{Hilbert}}$, in which the role of $g_{f}$ is played by $\wedge$.
@@
Lastly, recall that $f\in\mathsf{ext}_{pp}(\mathsf{Hilbert})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{Heyting}_{\to}$. Together with the left hand side of (41), this allows us to apply 10.5, obtaining that $\mathbb{Q}(\mathsf{Heyting}_{\to}[{L}_{f}])$ is a pp expansion of $\mathsf{Hilbert}$. By applying in succession the above display and the right hand side of (41), we obtain
@@
Now, let $f\in\mathsf{ext}_{pp}(\mathsf{PDL})$ be the binary implicit operation of $\mathsf{PDL}$ given by 8.25. Moreover, let $\to$ be a binary function symbol. Then the language ${L}_{f}={L}_{\mathsf{PDL}}\cup\{\to\}$ is an $f$-expansion of ${L}_{\mathsf{PDL}}$, in which the role of $g_{f}$ is played by $\to$.
@@
Finally, recall that $f\in\mathsf{ext}_{pp}(\mathsf{PDL})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{PDL}_{\textsc{si}}^{\textup{fg}}$. Together with the left hand side of (42), this allows us to apply 10.5, obtaining that $\mathbb{Q}(\mathsf{PDL}_{\textsc{si}}^{\textup{fg}}[{L}_{f}])$ is a pp expansion of $\mathsf{PDL}$. By applying in succession the above display and the right hand side of (42), we obtain
@@
The rest of this section is devoted to proving Theorems 10.3, 10.4, 10.5, 10.6, and 10.7. We postpone the proof of 10.3 and begin by proving 10.4.
@@
axiomatizes $\mathsf{K}[{L}_{\mathcal{F}}]$. To show that $\mathsf{M}=\mathsf{K}[{L}_{\mathcal{F}}]$ it is sufficient to prove that $\mathsf{K}[{L}_{\mathcal{F}}]$ is closed under $\mathbb{S}$. As $\mathsf{K}$ is a universal class, 2.1(iii) allows us to assume that $\Sigma$ consists of universal formulas. Together with the fact that each $\varphi_{f}(x_{1},\dots,x_{n},g_{f}(x_{1},\dots,x_{n}))$ is a conjunction of equations, this implies that $\mathsf{K}[{L}_{\mathcal{F}}]$ can be axiomatized by a set of universal formulas. Therefore, $\mathsf{K}[{L}_{\mathcal{F}}]$ is a universal class by 2.1(iii), and hence it is closed under $\mathbb{S}$. ∎
We are now ready to prove 10.3.
@@
As $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$ and $\mathsf{K}$ is closed under $\mathbb{O}$, we known that $\mathsf{K}$ is a universal class. Therefore, we can apply Theorem 8.4, obtaining that for each $\bm{A}\in\mathsf{K}$ there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. Since $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})\subseteq\mathsf{ext}(\mathsf{K})$ by assumption, for each $\bm{A}\in\mathsf{K}$ there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathcal{F}$. Together with the definition of $\mathsf{N}$, this yields $\mathsf{K}\subseteq\mathbb{S}(\mathsf{N})$. As $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$, we obtain $\mathsf{K}\subseteq\mathbb{O}(\mathsf{N})$. On the other hand, from $\mathsf{N}\subseteq\mathsf{K}$ and the assumption that $\mathsf{K}$ is closed under $\mathbb{O}$ it follows that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{K}$, whence $\mathsf{K}=\mathbb{O}(\mathsf{N})$. ∎
@@
Since the quasivariety $\mathsf{K}$ is closed under $\mathbb{I}\mathbb{S}\mathbb{P}$, the definition of $\mathsf{N}_{\textsc{rsi}}$ allows us to apply 9.10, obtaining $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathbb{S}((\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}))[{L}_{\mathcal{F}}])$. By Claim 10.21 this amounts to $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, we conclude that $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathsf{M}$.
@@
As $\bm{A}\in\mathsf{M}_{\textsc{rsi}}$, this implies that there exists $i\in I$ such that the map $p_{i}\circ h\colon\bm{A}\to p_{i}[h[\bm{A}]]$ is an isomorphism. Since $p_{i}[h[\bm{A}]]\leqslant\bm{A}_{i}[{L}_{\mathcal{F}}]$ and $\bm{A}_{i}\in\mathsf{N}_{\textsc{rsi}}$, we obtain $\bm{A}\in\mathbb{I}\mathbb{S}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])$. Lastly, applying in succession $\mathbb{I}\mathbb{S}=\mathbb{S}\mathbb{I}$, 9.8(i) for the case where $\mathbb{O}=\mathbb{I}$, the fact that $\mathsf{N}_{\textsc{rsi}}$ is closed under $\mathbb{I}$, and the inclusion $\mathsf{N}_{\textsc{rsi}}\subseteq\mathsf{K}_{\textsc{rsi}}$, we conclude that
@@
It remains to prove (iii). Assume that $\mathsf{K}$ is a variety and $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$. By 2.1(i) there exists a set of equations $\Sigma$ that axiomatizes $\mathsf{K}$. Then 10.4 yields that $\mathsf{M}$ is axiomatized by
@@
where $\varphi_{f}$ is the conjunction of equations defining $f\in\mathcal{F}$. Consequently, $\mathsf{M}$ is axiomatized by a set of equations. Thus, $\mathsf{M}$ is a variety by 2.1(i). ∎
@@
We detail only the case in which $\mathbb{O}=\mathbb{U}$, as the case in which $\mathbb{O}=\mathbb{Q}$ is handled analogously. Accordingly, assume that $\mathsf{K}=\mathbb{U}(\mathsf{N})$. As $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ by assumption, the class $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$. Therefore, it only remains to show that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])=\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])$. Since
@@
by assumption, we can apply 9.10, obtaining $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])=\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$. As $\mathsf{K}=\mathbb{U}(\mathsf{N})$, this amounts to $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. ∎
It only remains to prove 10.6. The proof hinges on the next observation, in which a nonconstant term is simply a term that is not a constant.
@@
    for every constant $c$ of $\mathsf{M}$ there exists a unary $f_{c}\in\mathsf{ext}_{eq}(\mathsf{K})$ such that $f_{c}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ is total and $c^{\bm{A}}=f_{c}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a)$ for all $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $a\in A$;
@@
    for every nonconstant term $t$ of $\mathsf{M}$ there exists $f_{t}\in\mathsf{ext}_{pp}(\mathsf{K})$ such that $t^{\bm{A}}=f_{t}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$;
@@
    for every $f\in\mathsf{imp}_{pp}(\mathsf{M})$ there exists $f_{*}\in\mathsf{imp}_{pp}(\mathsf{K})$ such that $f^{\bm{A}}=f_{*}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. Furthermore, if $f\in\mathsf{ext}_{pp}(\mathsf{M})$, then $f_{*}$ can be chosen in $\mathsf{ext}_{pp}(\mathsf{K})$.
@@
By assumption $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$. Therefore, ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Consequently, $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ is of the form ${L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$. This fact will be used repeatedly in the proof.
(i): Let $c$ be a constant of $\mathsf{M}$. Then $c$ belongs to ${L}_{\mathcal{F}}$, which is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. As $\mathcal{F}$-expansions are obtained by adding only functions symbols of positive arity, it follows that $c\in{L}_{\mathsf{K}}$. Then the equation $\varphi(x,y)=y\thickapprox c$ defines a unary $f_{c}\in\mathsf{ext}_{eq}(\mathsf{K})$ with the desired properties.
(ii): We proceed by induction on the construction of the nonconstant term $t(x_{1},\dots,x_{n})$. In the base case, $t=p(x_{1},\dots,x_{n})$ for a basic operation $p$ of $\mathsf{M}$. Therefore, $p\in{L}_{\mathcal{F}}$. As ${L}_{\mathcal{F}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$, we have two cases: either $p\in{L}_{\mathsf{K}}$ or there exists $h\in\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ such that $p=g_{h}$. In the first case, we let $f_{t}=\langle p^{\bm{A}}:\bm{A}\in\mathsf{K}\rangle$ and, in the second, $f_{t}=h$. In both cases, $f_{t}\in\mathsf{ext}_{pp}(\mathsf{K})$ and $t^{\bm{A}}=f_{t}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$.
In the inductive step, $t=p(t_{1},\dots,t_{m})$ for some $m$-ary $p\in{L}_{\mathcal{F}}$ and terms $t_{i}(x_{1},\dots,x_{n})$ of $\mathsf{M}$. By the inductive hypothesis there exist $f_{p},f_{t_{1}},\dots,f_{t_{m}}\in\mathsf{ext}_{pp}(\mathsf{K})$ satisfying the condition in the statement for $p,t_{1},\dots,t_{m}$, respectively. Define $f_{t}$ as the composition $f_{p}(f_{t_{1}},\dots,f_{t_{m}})$. As $f_{p},f_{t_{1}},\dots,f_{t_{m}}\in\mathsf{ext}_{pp}(\mathsf{K})$, from 8.7 it follows that $f_{t}\in\mathsf{ext}_{pp}(\mathsf{K})$ as well. Then consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $a_{1},\dots,a_{n}\in A$. By the inductive hypothesis $p^{\bm{A}}=f_{p}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ and $t_{i}^{\bm{A}}=f_{t_{i}}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $i\leqslant m$. Therefore, for each $i\leqslant m$,
@@
Recall that $\varphi$ defines the implicit operation $f$ of $\mathsf{M}$. Therefore, $\varphi$ is functional in $\mathsf{M}$ and, in particular, in $\mathsf{K}[{L}_{\mathcal{F}}]$. Together with Claim 10.23, this yields that $\psi$ is functional in $\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}$. As $\psi$ is equivalent to a pp formula by definition, we can apply 3.11, obtaining that $\psi$ defines some $f_{*}\in\mathsf{imp}_{pp}(\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}))$. Then recall that $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$ and, therefore, $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$. Consequently, we can apply Proposition 9.6, obtaining that $\mathsf{K}=\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\})$. Hence,
$f_{*}\in\mathsf{imp}_{pp}(\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}))=\mathsf{imp}_{pp}(\mathsf{K}).$

### v2

@@ 10. Primitive positive expansions
Let $f$ be the unary implicit operation of the quasivariety $\mathsf{RCRing}$ of reduced commutative rings given by Theorem 8.18 and recall from the same theorem that $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{RCRing})$. Moreover, let $(\,)^{*}$ be a unary function symbol. Then the language ${L}_{f}={L}_{\mathsf{RCRing}}\cup\{(\,)^{*}\}$ is an $f$-expansion of ${L}_{\mathsf{RCRing}}$, in which the role of $g_{f}$ is played by $(\,)^{*}$.
@@
where $\mathsf{Meadow}$ is the variety of meadows. Lastly, recall that $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{RCRing})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{Field}$. Together with the left hand side of the above display, this allows us to apply Theorem 10.5, obtaining that $\mathbb{Q}(\mathsf{Field}[{L}_{f}])$ is a pp expansion of $\mathsf{RCRing}$. By applying in succession the equality $\mathsf{Field}[{L}_{f}]=\mathsf{Field}^{*}$ and the right hand side of the above display, we obtain
@@
Now, let $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{Hilbert})$ be the binary implicit operation given by Theorem 8.22. Moreover, let $\wedge$ be a binary function symbol. Then the language ${L}_{f}={L}_{\mathsf{Hilbert}}\cup\{\wedge\}$ is an $f$-expansion of ${L}_{\mathsf{Hilbert}}$, in which the role of $g_{f}$ is played by $\wedge$.
@@
Lastly, recall that $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{Hilbert})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{Heyting}_{\to}$. Together with the left hand side of (41), this allows us to apply Theorem 10.5, obtaining that $\mathbb{Q}(\mathsf{Heyting}_{\to}[{L}_{f}])$ is a pp expansion of $\mathsf{Hilbert}$. By applying in succession the above display and the right hand side of (41), we obtain
@@
Now, let $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{PDL})$ be the binary implicit operation of $\mathsf{PDL}$ given by Theorem 8.25. Moreover, let $\to$ be a binary function symbol. Then the language ${L}_{f}={L}_{\mathsf{PDL}}\cup\{\to\}$ is an $f$-expansion of ${L}_{\mathsf{PDL}}$, in which the role of $g_{f}$ is played by $\to$.
@@
Finally, recall that $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{PDL})$ and that $f^{\bm{A}}$ is total for each $\bm{A}\in\mathsf{PDL}_{\textsc{si}}^{\textup{fg}}$. Together with the left hand side of (42), this allows us to apply Theorem 10.5, obtaining that $\mathbb{Q}(\mathsf{PDL}_{\textsc{si}}^{\textup{fg}}[{L}_{f}])$ is a pp expansion of $\mathsf{PDL}$. By applying in succession the above display and the right hand side of (42), we obtain
@@
The rest of this section is devoted to proving Theorems 10.3, 10.4, 10.5, 10.6, and 10.7. We postpone the proof of Theorem 10.3 and begin by proving Theorem 10.4.
@@
axiomatizes $\mathsf{K}[{L}_{\mathcal{F}}]$. To show that $\mathsf{M}=\mathsf{K}[{L}_{\mathcal{F}}]$ it is sufficient to prove that $\mathsf{K}[{L}_{\mathcal{F}}]$ is closed under $\mathbb{S}$. As $\mathsf{K}$ is a universal class, Theorem 2.1(iii) allows us to assume that $\Sigma$ consists of universal formulas. Together with the fact that each $\varphi_{f}(x_{1},\dots,x_{n},g_{f}(x_{1},\dots,x_{n}))$ is a conjunction of equations, this implies that $\mathsf{K}[{L}_{\mathcal{F}}]$ can be axiomatized by a set of universal formulas. Therefore, $\mathsf{K}[{L}_{\mathcal{F}}]$ is a universal class by Theorem 2.1(iii), and hence it is closed under $\mathbb{S}$. ∎
We are now ready to prove Theorem 10.3.
@@
As $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$ and $\mathsf{K}$ is closed under $\mathbb{O}$, we known that $\mathsf{K}$ is a universal class. Therefore, we can apply Theorem 8.4, obtaining that for each $\bm{A}\in\mathsf{K}$ there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathsf{ext}(\mathsf{K})$. Since $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})\subseteq\mathsf{ext}(\mathsf{K})$ by assumption, for each $\bm{A}\in\mathsf{K}$ there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $f^{\bm{B}}$ is total for each $f\in\mathcal{F}$. Together with the definition of $\mathsf{N}$, this yields $\mathsf{K}\subseteq\mathbb{S}(\mathsf{N})$. As $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$, we obtain $\mathsf{K}\subseteq\mathbb{O}(\mathsf{N})$. On the other hand, from $\mathsf{N}\subseteq\mathsf{K}$ and the assumption that $\mathsf{K}$ is closed under $\mathbb{O}$ it follows that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{K}$, whence $\mathsf{K}=\mathbb{O}(\mathsf{N})$. ∎
@@
Since the quasivariety $\mathsf{K}$ is closed under $\mathbb{I}\mathbb{S}\mathbb{P}$, the definition of $\mathsf{N}_{\textsc{rsi}}$ allows us to apply Proposition 9.10, obtaining $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathbb{S}((\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}))[{L}_{\mathcal{F}}])$. By Claim 10.21 this amounts to $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, we conclude that $\mathbb{I}\mathbb{S}\mathbb{P}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])=\mathsf{M}$.
@@
As $\bm{A}\in\mathsf{M}_{\textsc{rsi}}$, this implies that there exists $i\in I$ such that the map $p_{i}\circ h\colon\bm{A}\to p_{i}[h[\bm{A}]]$ is an isomorphism. Since $p_{i}[h[\bm{A}]]\leqslant\bm{A}_{i}[{L}_{\mathcal{F}}]$ and $\bm{A}_{i}\in\mathsf{N}_{\textsc{rsi}}$, we obtain $\bm{A}\in\mathbb{I}\mathbb{S}(\mathsf{N}_{\textsc{rsi}}[{L}_{\mathcal{F}}])$. Lastly, applying in succession $\mathbb{I}\mathbb{S}=\mathbb{S}\mathbb{I}$, Proposition 9.8(i) for the case where $\mathbb{O}=\mathbb{I}$, the fact that $\mathsf{N}_{\textsc{rsi}}$ is closed under $\mathbb{I}$, and the inclusion $\mathsf{N}_{\textsc{rsi}}\subseteq\mathsf{K}_{\textsc{rsi}}$, we conclude that
@@
It remains to prove (iii). Assume that $\mathsf{K}$ is a variety and $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$. By Theorem 2.1(i) there exists a set of equations $\Sigma$ that axiomatizes $\mathsf{K}$. Then Theorem 10.4 yields that $\mathsf{M}$ is axiomatized by
@@
where $\varphi_{f}$ is the conjunction of equations defining $f\in\mathcal{F}$. Consequently, $\mathsf{M}$ is axiomatized by a set of equations. Thus, $\mathsf{M}$ is a variety by Theorem 2.1(i). ∎
@@
We detail only the case in which $\mathbb{O}=\mathbb{U}$, as the case in which $\mathbb{O}=\mathbb{Q}$ is handled analogously. Accordingly, assume that $\mathsf{K}=\mathbb{U}(\mathsf{N})$. As $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ by assumption, the class $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$. Therefore, it only remains to show that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])=\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])$. Since
@@
by assumption, we can apply Proposition 9.10, obtaining $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])=\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$. As $\mathsf{K}=\mathbb{U}(\mathsf{N})$, this amounts to $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. ∎
It only remains to prove Theorem 10.6. The proof hinges on the next observation, in which a nonconstant term is simply a term that is not a constant.
@@
    for every constant $c$ of $\mathsf{M}$ there exists a unary $f_{c}\in\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ such that $f_{c}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ is total and $c^{\bm{A}}=f_{c}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a)$ for all $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $a\in A$;
@@
    for every nonconstant term $t$ of $\mathsf{M}$ there exists $f_{t}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ such that $t^{\bm{A}}=f_{t}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$;
@@
    for every $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{M})$ there exists $f_{*}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ such that $f^{\bm{A}}=f_{*}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. Furthermore, if $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{M})$, then $f_{*}$ can be chosen in $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$.
@@
By assumption $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$. Therefore, ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Consequently, $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ is of the form ${L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$. This fact will be used repeatedly in the proof.
(i): Let $c$ be a constant of $\mathsf{M}$. Then $c$ belongs to ${L}_{\mathcal{F}}$, which is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. As $\mathcal{F}$-expansions are obtained by adding only functions symbols of positive arity, it follows that $c\in{L}_{\mathsf{K}}$. Then the equation $\varphi(x,y)=y\thickapprox c$ defines a unary $f_{c}\in\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ with the desired properties.
(ii): We proceed by induction on the construction of the nonconstant term $t(x_{1},\dots,x_{n})$. In the base case, $t=p(x_{1},\dots,x_{n})$ for a basic operation $p$ of $\mathsf{M}$. Therefore, $p\in{L}_{\mathcal{F}}$. As ${L}_{\mathcal{F}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$, we have two cases: either $p\in{L}_{\mathsf{K}}$ or there exists $h\in\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ such that $p=g_{h}$. In the first case, we let $f_{t}=\langle p^{\bm{A}}:\bm{A}\in\mathsf{K}\rangle$ and, in the second, $f_{t}=h$. In both cases, $f_{t}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and $t^{\bm{A}}=f_{t}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$.
In the inductive step, $t=p(t_{1},\dots,t_{m})$ for some $m$-ary $p\in{L}_{\mathcal{F}}$ and terms $t_{i}(x_{1},\dots,x_{n})$ of $\mathsf{M}$. By the inductive hypothesis there exist $f_{p},f_{t_{1}},\dots,f_{t_{m}}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ satisfying the condition in the statement for $p,t_{1},\dots,t_{m}$, respectively. Define $f_{t}$ as the composition $f_{p}(f_{t_{1}},\dots,f_{t_{m}})$. As $f_{p},f_{t_{1}},\dots,f_{t_{m}}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, from Corollary 8.7 it follows that $f_{t}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ as well. Then consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $a_{1},\dots,a_{n}\in A$. By the inductive hypothesis $p^{\bm{A}}=f_{p}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ and $t_{i}^{\bm{A}}=f_{t_{i}}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $i\leqslant m$. Therefore, for each $i\leqslant m$,
@@
Recall that $\varphi$ defines the implicit operation $f$ of $\mathsf{M}$. Therefore, $\varphi$ is functional in $\mathsf{M}$ and, in particular, in $\mathsf{K}[{L}_{\mathcal{F}}]$. Together with Claim 10.23, this yields that $\psi$ is functional in $\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}$. As $\psi$ is equivalent to a pp formula by definition, we can apply Corollary 3.11, obtaining that $\psi$ defines some $f_{*}\in\mathsf{imp}_{\textsc{pp}}(\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}))$. Then recall that $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$ and, therefore, $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Consequently, we can apply Proposition 9.6, obtaining that $\mathsf{K}=\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\})$. Hence,
$f_{*}\in\mathsf{imp}_{\textsc{pp}}(\mathbb{S}(\{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}:\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]\}))=\mathsf{imp}_{\textsc{pp}}(\mathsf{K}).$